OpenStudy (anonymous):

Your an engineer working on Ares 1, NASA's replacement for the space shuttles, performance specs call for a first -stage rocket capable of accelerating a total mass of 660Mg vertically from rest to 7700 km/h in 1.7 min. Your asked to determine the required thrust (force). What do you report?

OpenStudy (shane_b):

Ok, first I'd convert everything to the same units: 7700 km/h = 2138.89 m/s 660Mg = 660000kg 1.7min = 102s Ok, let's see what our acceleration needs to be to reach that speed. Using one of the kinematic equations you get: \[V_f=V_i+at\]\[2138.89m/s=0m/s+a_{net}(102s)\]\[a_{net}=20.97m/s^2\] Now let's not forget to add in the acceleration we need to overcome the acceleration due to gravity:\[a_{total}=a_{net}+a_g=20.97m/s^2 + 9.81m/s^2=30.78m/s^2\] Ok finally, let's calculate the force needed now that we know the total acceleration required:\[F=ma=(660000kg)(30.78m/s^2)=20314800N=20.3148 MN\]